# A search for Galactic post-AGB stars in Gaia DR3††thanks: Tables A6 and A7 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A

### v1

@@ Abstract
Context. When low and intermediate-mass stars leave the Asymptotic Giant Branch (AGB) phase, and before they reach the Planetary Nebulae (PN) stage, they enter a very brief and rather puzzling stellar evolutionary stage named post-AGB. The post-AGB phase has a very short duration, of the order of a few thousand years at most, so the number of objects confirmed in this phase is really small, and thus, our understanding of this elusive stellar evolutionary stage is very constrained.
Aims. To provide a reliable catalogue of galactic post-AGB stars together with their physical and evolutionary properties obtained through Gaia DR3 astrometry and photometry. As an added product, we provide information for a sample of other types of stellar objects, whose observational properties mimic those of post-AGB stars.
Methods. Post-AGB stars are characterised by their infrared excesses and high luminosities. The publication of precise parallaxes in Gaia DR3 made it possible to calculate accurate distances and to revise the derivation of luminosities for post-AGB candidates to discard objects outside the expected luminosity range. We started by identifying post-AGB stars or possible candidates from the bibliography and, then, searched for their Gaia DR3 counterpart sources. Using the available photometry, interstellar extinction, literature spectroscopically derived temperatures or spectral types and parallax-derived distances, we fitted their Spectral Energy Distributions (SED) and we estimated their luminosities and circumstellar extinctions. When compared to models, luminosity values allowed us to disclose objects that are likely post-AGB stars from other target types. Their position on the Hertzsprung-Russell (HR) diagram allows direct comparison with updated post-AGB evolutionary tracks and an estimation of their masses and evolutionary ages.
Results. We obtained a sample of 69 reliable post-AGB candidates that meet our classification criteria, providing their coordinates, distances, effective temperature, interstellar and circumstellar extinction, luminosity, mass, and evolutionary age. In addition, similar data for other stellar objects in our initial compilation, such as supergiant stars and Young Stellar Objects (YSOs), is provided. Our identifications and parameters are compared with others found in recent literature about the subject.
Conclusions. We have filtered out the data that have the best precision in parallaxes and distances to obtain more accurate luminosities, which consequently, allows us to classify with confidence the objects of the sample among different stellar phases. In turn, this allows us to provide a small but reliable sample of post-AGB objects. Derived mean evolutionary time and average mass values are in agreement with theoretical expectations, and with the mass mean value obtained in a previous work for the subsequent evolutionary stage, the Planetary Nebula stage.
@@
The stellar phase known as post Asymptotic Giant Branch (post-AGB) is a very fast (a few thousand years) and quite unknown phase, that takes place at the end of the low and intermediate-mass star lifetime, after the AGB phase and before the Planetary Nebula (PN) stage, where the star ionises the previously ejected envelope. There is no exact definition for the beginning of the post-AGB phase, and its onset depends on the stellar mass and metal content. Furthermore, the departure from the AGB phase is defined somewhat arbitrarily in the stellar evolution models. For Vassiliadis & Wood (1993) the AGB mass-loss is terminated when the envelope mass is reduced to such a value that the stellar effective temperature increases beyond a reference value by an amount of log(Teff)=0.3. Other authors have different criteria. In Miller Bertolami (2016), they set the onset when, due to the stellar winds, the H-rich envelope mass drops below 1% of the stellar mass.
@@
During the post-AGB phase, the star will evolve at almost constant luminosity towards hotter effective temperatures, while its envelope will expand into the interstellar medium. Due to the high stellar temperature, this envelope composed of gas and dust will start to be ionised (see e.g. Villaver et al. 2002). At this moment, the star will enter the proto-planetary nebulae phase. As in González-Santamaría et al. (2021) we can assume a minimum stellar temperature of 13 000 K for a transition stage from a preplanetary nebula (Weidmann et al. 2020) to about 24 000 K for a complete ionisation of the nebula (Kwok, 2000).
@@
Over the last years, several studies have been carried out to identify post-AGB stars to better characterise this brief stellar phase. Garcia-Lario et al. (1997), based on their IRAS infrared colours, identified 110 possible post-AGBs, most of them without optical counterparts. At the beginning of this century, Suárez et al. (2006) provided optical counterparts for more than 100 post-AGB candidates, previously proposed as such based on their IRAS fluxes. Almost in parallel, a more extensive catalogue of post-AGB candidates was presented by Szczerba et al. (2007), which was later extended in Szczerba et al. (2012). This is known as the Torun catalogue of post-AGB stars, which is the largest one to date containing 296 sources classified as either Likely or Possible post-AGBs. The sample of post-AGB stars has been expanded to the Large Magellanic Cloud (LMC) by Kamath et al. (2015), including also post-RGB (Red Giant Branch) stars. These stars are believed to be produced from binary interactions in the RGB phase, and their observational properties mimic those of post-AGB stars, although they are not expected to reach luminosities as high as those of post-AGB stars (Kamath et al. 2015). In a previous article (González-Santamaría et al. 2021), we found that a fraction close to 50% of the central stars of PNe are red, and therefore possibly unresolved binary systems. Finding the relative number of stars of one type and another, single or binary, at those evolutionary stages, can give us clues about the fraction of binary stars with evolved companions.
@@
Recent work by Kamath et al. (2022) studies the evolutionary state of 31 galactic post-AGB candidates with chemical abundance information. By using Gaia EDR3 distances and known photometry, they build the Spectral Energy Distribution (SED) for these objects, and by fitting them to models, they estimate their luminosities and temperatures. It is important to note that 20 of the objects analysed have quite poor astrometry in Gaia DR3.
@@
In Parthasarathy et al. (2020), the properties of 8 post-AGB candidates were analysed making use of Gaia DR2 astrometric and photometric data. Also using Gaia, in this case DR3, Oudmaijer et al. (2022) investigate the nature as possible post-AGB of 249 objects. It is noteworthy that most of the objects here selected present important uncertainties in their parallaxes, which leads to very unreliable distance values. This work considers good parallaxes those with relative error () between 10% and 100%. Additionally, we detected some inconsistencies in this study that we shall comment on in section 3.
@@
We aim to go further into the identification and analysis of the properties of bona fide post-AGB stars by making a selection of objects with accurate astrometric measurements in Gaia DR3, thus allowing a precise location in the Hertzsprung-Russell (HR) diagram and, consequently, a quite reliable classification as post-AGB objects. Selecting among Galactic post-AGB stars candidates with good astrometry in Gaia DR3, in a first approximation, would exclude most astrometric binaries from it. The general rule is that the threshold of Gaia astrometric quality parameter Renormalised Unit Weight Error (RUWE) 1.4 is used to indicate single well-behaved solutions (Lindegren et al. 2018, 2021). Sources whose observations are inconsistent with Gaia astrometric 5-parameter model could be caused by binarity (Lindegren et al. 2018) or other factors that cause the photocentre of the source to wobble during the Gaia observation window. In summary, a restrictive criterion in the astrometric quality will help us to select objects with a higher probability of being individual sources, and on the other hand, it will help us to better estimate their luminosity and evolutionary stage. The incidence of binaries in our resulting sample, however, will be addressed by studying the SED morphology and also in the literature context.
@@
We started by collecting a sample of post-AGB candidates as complete as possible from the currently available catalogues (Sect. 2) and we implemented restrictive filtering over the astrometric quality of the objects, to only keep those with the most accurate distance values. Concretely, we gathered a general sample of 964 post-AGB candidates from the literature of which we filtered out a subset of 178 objects with accurate astrometric measurements in Gaia DR3. A good distance determination is not enough to obtain a reliable adjustment of the SED; it is necessary to have additional information about the temperature and/or extinction in the direction of the source, allowing to consistently derive the luminosity of the object. To better constrain the value of the total extinction, interstellar and circumstellar, in the direction of the source, we opted to limit our work to those objects with available interstellar extinction measures in the literature. From the previous sample of 178 objects, we kept 146 galactic post-AGB candidate stars with literature values of their interstellar extinction. In this last sample, the information available in the Simbad database as well as images of every sky field in Aladin Sky Atlas (Bonnarel et al. 2000) were analysed. 28 sources were finally discarded for further analysis because either they have been already classified as PNe (5 objects), or they have effective temperatures in the literature higher than 24 000 K (11 objects), or the identification of the optical counterpart is dubious (12 objects). Our final working set turned out to be 118 objects. We found that for about 67% of them, temperature values were derived from spectral analysis, with a fraction of them coming from high-resolution spectral analysis. For the rest, average temperatures corresponding to their spectral types were used. Section 3.2 describes in detail the problems associated with the different quality of temperature determinations used in our work.
@@
Photometry in a wide spectral region compiled by the spanish Virtual Observatory SED Analyser (VOSA111http://svo2.cab.inta-csic.es/theory/vosa) alowed us to build the SED for each object. The knowledge of distances and temperatures allowed us to obtain luminosities and to estimate the total extinction from the SED fitting (Sect. 3.3). We have used current knowledge of interstellar extinction in the direction of every object to verify the consistency of the total extinction obtained in the fits.
@@
Based on evolutionary tracks, we used the distribution of objects in an HR diagram to confirm 69 objects out of 118 candidates as post-AGB stars. Meanwhile, some other objects could be classified as Horizontal Branch stars (3 objects), luminous supergiants (3 objects), Young Stellar Objects (YSO –14 candidates), and the rest would remain as unconfirmed post-AGB candidates, –details to be given in Sec. 3.5. In Section 3.6, this classification is compared with those presented in other recent papers about the subject. In Sect. 4, we focussed on analysing the evolutionary properties of the set of 69 objects that we identify as belonging to the post-AGB stage. In Sect. 5, we comment on some interesting objects, and, finally, in Sect. 6, we summarise our conclusions.
2 Sample selection: methodology
@@
The first step in this research was to collect all the objects from the literature catalogued as confirmed or possible, post-AGB stars. For this purpose, we used the online222https://fox.ncac.torun.pl/camkweb/postagb2.php Torun catalogue of Galactic post-AGB stars (Szczerba et al. 2012), the Simbad astronomical database and the spectroscopic atlas of post-AGB and planetary nebulae by Suárez et al. (2006) (from now on Suarez et al. catalogue).
@@
Concretely, we gathered all objects catalogued as Likely (209) or Possible (87) in the Torun catalogue, as Post-AGB (331) or post-AGB Candidate (507) in the Simbad database and as Post-AGB (102) in the Suarez et al. catalogue. To obtain the union of the three catalogues, firstly, we matched the Torun objects with those in Simbad using a 1 arcsec cross-match radius. As a result, we obtained a set of 929 objects. Subsequently, we matched those with Suarez et al. catalogue (by the same method and using coordinates from Torun when possible), obtaining a final sample of 964 objects.
@@
The next step was to cross-match our list of objects with the Gaia DR3 archive to obtain their parallaxes, and consequently their distances. We used again a searching radius of 1 arcsec from the literature coordinates. Note that in some cases, the coordinates from Simbad or Suárez et al. catalogues differ slightly from the coordinates from the Torun catalogue; this can lead to discrepancies in the identification of the Gaia DR3 source. We decided to prioritise the Torun coordinates, as they come from a revised compilation of post-AGB stars. As a result, we were able to identify 843 objects as Gaia DR3 sources, around 87% of the whole sample. Finally, we discarded those objects with unknown distances in Bailer-Jones et al. (2021), and we ended up with 765 objects.
@@
In Gaia DR3, parallaxes () show a bias or zero point () that should be considered and subtracted from the measured value to obtain the actual parallax (). According to Lindegren et al. (2021) this zero point has a mean value of -17 as, although its value varies depending on the star celestial position, colour and magnitude. Gaia web provides a Python code to estimate this zero point333https://gitlab.com/icc-ub/public/Gaiadr3_zeropoint. After obtaining the zero points, we corrected the parallaxes with the following simple expression:
@@
Uncertainties in Gaia parallaxes include both the value given in Gaia DR3 archive, the internal uncertainty, as well as a systematic uncertainty that depends on the source brightness. We followed the prescriptions given in Fabricius et al. (2021) to estimate the parallax total uncertainties of our objects.
@@
From the inverse of the parallaxes, it is possible to estimate the stellar distances, but this simple approach is only valid for objects with relative uncertainties as low as 10%. In general, it is not our case, so we used the distances calculated by a Bayesian approach for the Milky Way stars by Bailer-Jones et al. (2021), which consists of assuming an a priori probability volume density of stars in the Galaxy that decreases exponentially on an appropriate distance scale. This method not only provides an estimated distance for a source but also gives low and high error distance bounds.
@@
To infer useful stellar properties that depend on distance, such as luminosity, it is important to have precise distances to these objects. So we decided to apply filtering to our sample according to the distance uncertainties as well as to the astrometric quality. Astrometric quality indices such as UWE (Unit Weight Error, see Gaia official website) and RUWE (Renormalised Unit Weight Error) beyond certain boundaries prevent Gaia users from errors in the astrometric solution that can be due to binarity or to irregularities in the fitted source. We decided to use the same filtering criteria as that used in one of our previous works, where we did a similar study but over the galactic sample of planetary nebulae central stars (González-Santamaría et al. 2021). This filtering consists of the following constraints: relative error in parallax and distance (for both lower and upper bounds) below 30% and astrometric quality parameters UWE and RUWE below certain threshold values recommended in Gaia documentation (UWE¡1.96 or RUWE¡1.4). After applying these constraints, we ended up with a set of 178 objects that we consider to have good astrometric measurements in Gaia DR3.
@@
If we analyse the galactic distribution of these stars we appreciate a small cluster of 21 objects in a region of mid-south latitudes (see Fig. 1) at longitudes coincident with those of the direction of the LMC (Large Magellanic Cloud). Furthermore, we verified that all of these objects are included in the van Aarle et al. (2011) catalogue of LMC post-AGB stars, while 10 of them are also included in Kamath et al. (2015). So, we decided to exclude those stars from our present study and focus on the remaining 157 Galactic post-AGB candidates.
@@
In Table LABEL:table:candidates we provide general data of the objects in this later sample: a running number, name, Gaia DR3 ID, J2000 coordinates, G magnitude, interstellar extinction in the V-band, spectral type, reference for identification as post-AGB (Torun, Simbad and Suárez et al. catalogues), and a flag providing information about binarity, variability and the reason for excluding the source for further analysis, as it was mentioned in the previous section, together with a reference.
@@
The distance distribution of all these post-AGB candidates is shown in Fig. 2. We will not attempt to do any analysis of the completeness of the sample since it presents several observational biases, among which is the impossibility of detecting the optical counterpart of some of the infrared sources with data in the IRAS catalogue.
3 Galactic post-AGB stars identification

### v2

# A search for Galactic post-asymptotic giant branch stars in Gaia DR3††thanks: Tables A6 and A7 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A

@@ Abstract
Context. When low- and intermediate-mass stars leave the asymptotic giant branch (AGB) phase, and before they reach the planetary nebula stage, they enter a very brief and rather puzzling stellar evolutionary stage called post-AGB stage. The post-AGB phase lasts very briefly, about a few thousand years at most. The number of objects that are confirmed in this phase therefore is really small, and our understanding of this elusive stellar evolutionary stage is accordingly very limited.
Aims. We provide a reliable catalogue of Galactic post-AGB stars together with their physical and evolutionary properties obtained through Gaia DR3 astrometry and photometry. As an added product, we provide information for a sample of other types of stellar objects, whose observational properties mimic those of post-AGB stars.
Methods. Post-AGB stars are characterised by their infrared excesses and high luminosities. The publication of precise parallaxes in Gaia DR3 made it possible to calculate accurate distances and to revise the derivation of luminosities for post-AGB candidates, so that objects outside the expected luminosity range can be discarded. We started by identifying post-AGB stars or possible candidates from the bibliography, and we then searched for their Gaia DR3 counterpart sources. Using the available photometry, interstellar extinction, spectroscopically derived temperatures or spectral types and parallax-derived distances from the literature, we fitted their spectral energy distributions and estimated their luminosities and circumstellar extinctions. By a comparison to models, the luminosity values allowed us to determine which objects are likely post-AGB stars from other target types. Their position in the Hertzsprung-Russell diagram allows a direct comparison with updated post-AGB evolutionary tracks and an estimation of their masses and evolutionary ages.
Results. We obtained a sample of 69 reliable post-AGB candidates that meet our classification criteria, which provide their coordinates, distances, effective temperature, interstellar and circumstellar extinction, luminosity, mass, and evolutionary age. In addition, we provide similar data for other stellar objects in our initial compilation, such as supergiant stars and young stellar objects. Our identifications and parameters are compared with others found in the recent literature for the subject.
Conclusions. We selected the data with the best precision in parallax and distance to obtain more accurate luminosities, which allowed us to confidently classify the objects of the sample in different stellar phases. In turn, this allowed us to provide a small but reliable sample of post-AGB objects. The derived mean evolutionary time and average mass values agree with theoretical expectations and with the mean mass value obtained in a previous work for the subsequent evolutionary stage, the planetary nebula stage.
@@
The stellar phase known as post-asymptotic giant branch (post-AGB) stage is a very fast (a few thousand years) and quite unknown phase that takes place at the end of the lifetimes of low- and intermediate-mass stars, after the AGB phase and before the planetary nebula stage, in which the star ionises the previously ejected envelope. The beginning of the post-AGB phase is not exactyl determined, and its onset depends on the stellar mass and metal content. Furthermore, the departure from the AGB phase is defined somewhat arbitrarily in the stellar evolution models. For Vassiliadis & Wood (1993) the AGB mass loss terminates when the envelope mass is reduced to a value at which the stellar effective temperature increases beyond a reference value by an amount of log(Teff)=0.3. Other authors have different criteria. Miller Bertolami (2016) set the onset at the point in time when as a result of stellar winds, the H-rich envelope mass drops below 1% of the stellar mass.
@@
During the post-AGB phase, the star evolves at an almost constant luminosity towards hotter effective temperatures, while its envelope expands into the interstellar medium. Due to the high stellar temperature, this envelope composed of gas and dust begins to be ionised (see e.g. Villaver et al. 2002). At this moment, the star enters the protoplanetary nebula phase. As in González-Santamaría et al. (2021), we can assume a minimum stellar temperature of 13 000 K for a transition stage from a preplanetary nebula (Weidmann et al. 2020) to about 24 000 K for a complete ionisation of the nebula (Kwok, 2000).
@@
Several studies have recently been carried out to identify post-AGB stars with the aim to better characterise this brief stellar phase. Garcia-Lario et al. (1997) identified 110 possible post-AGBs based on their IRAS111InfraRed Astronomical Satellite infrared colours, most of which lack an optical counterparts. At the beginning of this century, Suárez et al. (2006) provided optical counterparts for more than 100 post-AGB candidates that were previously proposed as such based on their IRAS fluxes. Almost in parallel, a more extensive catalogue of post-AGB candidates was presented by Szczerba et al. (2007), which was later extended in Szczerba et al. (2012). This is known as the Torun catalogue of post-AGB stars and is the largest catalgoue to date. It contains 296 sources classified as either likely or possible post-AGBs. The sample of post-AGB stars has been expanded to the Large Magellanic Cloud (LMC) by Kamath et al. (2015). They also included post-red giant branch (RGB) stars. These stars are thought to be produced from binary interactions in the RGB phase, and their observational properties mimic those of post-AGB stars, although they are not expected to reach luminosities as high as those of post-AGB stars (Kamath et al. 2015). In a previous article (González-Santamaría et al. 2021), we found that a fraction close to 50% of the central stars of PNe are red and might therefore be unresolved binary systems. Finding the relative number of stars of one type and of single or binary type at these evolutionary stages can give us clues about the fraction of binary stars that has evolved companions.
@@
Kamath et al. (2022) recently studied the evolutionary state of 31 Galactic post-AGB candidates with chemical abundance information. By using Gaia EDR3 distances and known photometry, they built the spectral energy distribution (SED) for these objects, and by fitting them to models, they estimated their luminosities and temperatures. It is important to note that 20 of the objects they analysed have quite poor astrometry in Gaia DR3.
@@
Parthasarathy et al. (2020) analysed the properties of 8 post-AGB candidates based on Gaia DR2 astrometric and photometric data. Also using Gaia, in this case, DR3, Oudmaijer et al. (2022) investigated their nature as possible post-AGB of 249 objects. It is noteworthy that most of the objects they selected have large uncertainties in their parallaxes, which leads to very unreliable distance values. They considered good parallaxes to have a relative error () between 10% and 100%. Additionally, we detected some inconsistencies in their study on which we comment in section 3.
@@
We aim to identify and analyse the properties of bona fide post-AGB stars in more detail by selecting objects with accurate astrometric measurements in Gaia DR3, which allows us to precisely locate them in the Hertzsprung-Russell (HR) diagram. This consequently leads to a quite reliable classification of these stars as post-AGB objects. If we were to select objects from Galactic post-AGB stars candidates with good astrometry in Gaia DR3, we would exclude in a first approximation most astrometric binaries. The general rule is that the threshold of the Gaia astrometric quality parameter called renormalised unit weight error (RUWE) 1.4 is used to indicate single well-behaved solutions (Lindegren et al. 2018, 2021). The inconsistency of source observations with the Gaia astrometric five-parameter model could be caused by binarity (Lindegren et al. 2018) or other factors that cause the photocentre of the source to wobble during the Gaia observation window. In summary, a restrictive criterion in the astrometric quality helps us to select objects that are more likely to be individual sources, and on the other hand, it helps us to estimate their luminosity and evolutionary stage better. The incidence of binaries in our resulting sample is addressed by studying the SED morphology, and we also consider this in the context of the literature.
@@
We started by collecting a sample of post-AGB candidates that was as complete as possible from the currently available catalogues (Sect. 2), and we implemented restrictive filtering over the astrometric quality of the objects so that we only kept those with the most accurate distance values. We gathered a general sample of 964 post-AGB candidates from the literature, out of which we filtered a subset of 178 objects with accurate astrometric measurements in Gaia DR3. A good distance determination is not enough to obtain a reliable adjustment of the SED; additional information about the temperature and/or extinction in the direction of the source is required for a consistently derived luminosity of the object. To better constrain the value of the total extinction, interstellar and circumstellar, in the direction of the source, we opted to limit our work to objects with available interstellar extinction measurements in the literature. Of the previous sample of 178 objects, we kept 146 Galactic post-AGB candidate stars with literature values of their interstellar extinction. In this last sample, the information available in the Simbad database as well as images of every sky field in the Aladin Sky Atlas (Bonnarel et al. 2000) were analysed. We finally discarded 28 sources from the further analysis because they were either already classified as PNe (5 objects), had higher effective temperatures in the literature than 24 000 K (11 objects), or because the identification of the optical counterpart was dubious (12 objects). Our final working set was 118 objects. We found that the temperatures for about 67% of them were derived from spectral analysis, and some of these from a high-resolution spectral analysis. The average temperatures for the remainder corresponding to their spectral types were used. Section 3.2 describes the problems associated with the different quality of the temperature determinations we used in detail.
@@
Photometry in a wide spectral region compiled by the Spanish Virtual Observatory SED Analyser (VOSA222http://svo2.cab.inta-csic.es/theory/vosa/) allowed us to build the SED for each object. The knowledge of distances and temperatures allowed us to obtain luminosities and to estimate the total extinction from the SED fitting (Sect. 3.3). We used the current knowledge of interstellar extinction in the direction of every object to verify the consistency of the total extinction we obtained in the fits.
@@
Based on evolutionary tracks, we used the distribution of objects in an HR diagram to confirm 69 objects out of 118 candidates as post-AGB stars. Some other objects could be classified as horizontal branch stars (3 objects), luminous supergiants (3 objects), young stellar objects (YSOs; 14 candidates), and the remaining objects are unconfirmed post-AGB candidates. The details are given in Sec. 3.5. In Section 3.6, this classification is compared with the classifications presented in other recent papers about the subject. In Sect. 4, we focus on analysing the evolutionary properties of the set of 69 objects that we identify as belonging to the post-AGB stage. In Sect. 5, we comment on some interesting objects, and in Sect. 6, we summarise our conclusions.
2 Sample selection: Method
@@
The first step in this research was to collect all the objects from the literature that were catalogued as confirmed or possible post-AGB stars. For this purpose, we used the online333https://fox.ncac.torun.pl/camkweb/postagb2.php Torun catalogue of Galactic post-AGB stars (Szczerba et al. 2012), the Simbad astronomical database, and the spectroscopic atlas of post-AGB and planetary nebulae by Suárez et al. (2006) (from now, the on Suarez et al. catalogue).
@@
We gathered all objects that were catalogued as likely (209) or possible (87) in the Torun catalogue, as post-AGB (331) or post-AGB candidate (507) in the Simbad database, and as post-AGB (102) in the Suarez et al. catalogue. To combine the three catalogues, we first matched the Torun objects with those in Simbad using a cross-match radius of one arcsecond. As a result, we obtained a set of 929 objects. Subsequently, we matched them with Suarez et al. catalogue (by the same method and using coordinates from Torun when possible), obtaining a final sample of 964 objects.
@@
The next step was to cross-match our list of objects with the Gaia DR3 archive to obtain their parallaxes and distances. We again used a search radius of 1 arcsecond from the literature coordinates. In some cases, the coordinates from Simbad or from the Suárez et al. catalogues differ slightly from the coordinates from the Torun catalogue. This can lead to discrepancies in the identification of the Gaia DR3 source. We decided to prioritise the Torun coordinates because they come from a revised compilation of post-AGB stars. As a result, we were able to identify 843 objects as Gaia DR3 sources. This is about 87% of the whole sample. Finally, we discarded objects with unknown distances in Bailer-Jones et al. (2021), which left 765 objects.
@@
In Gaia DR3, the parallaxes () show a bias or zero point () that should be considered and subtracted from the measured value to obtain the actual parallax (). According to Lindegren et al. (2021), this zero point has a mean value of -17 as, although its value varies depending on the celestial position, colour, and magnitude of the star. The Gaia web provides a Python code for estimating this zero point 444https://gitlab.com/icc-ub/public/gaiadr3_zeropoint. After obtaining the zero points, we corrected the parallaxes with the following simple expression:
@@
The uncertainties in the Gaia parallaxes include the value given in the Gaia DR3 archive, the internal uncertainty, and a systematic uncertainty that depends on the source brightness. We followed the prescriptions given in Fabricius et al. (2021) to estimate the total parallax uncertainties of our objects.
@@
From the inverse of the parallaxes, it is possible to estimate the stellar distances, but this simple approach is only valid for objects with relative uncertainties as low as 10%. In general, this is not the case for our data, and we therefore used the distances calculated by a Bayesian approach for Milky Way stars by Bailer-Jones et al. (2021), which consists of assuming an a priori probability volume density of stars in the Galaxy that decreases exponentially on an appropriate distance scale. This method not only provides an estimated distance for a source, but also gives low and high error distance bounds.
@@
To infer useful stellar properties that depend on distance, such as luminosity, it is important to have precise distances to these objects. We therefore decided to apply filtering to our sample according to the distance uncertainties and the astrometric quality. Astrometric quality indices such as the unit weight error (UWE; see the official Gaia website) and RUWE beyond certain boundaries prevent Gaia users from errors in the astrometric solution that can be due to binarity or to irregularities in the fitted source. We decided to use the same filtering criteria as we used in one of our previous works, where we conducted a similar study, but on the Galactic sample of central stars in planetary nebulae (González-Santamaría et al. 2021). This filtering consisted of the following constraints: The relative error in parallax and distance (for the lower and upper bounds) below 30% and the astrometric quality parameters UWE and RUWE below certain threshold values that are recommended in the Gaia documentation (UWE¡1.96 or RUWE¡1.4). After applying these constraints, the set contained 178 objects that we consider to have good astrometric measurements in Gaia DR3.
@@
The Galactic distribution of these stars shows a small cluster of 21 objects in a region of mid-south latitudes (see Fig. 1) at longitudes coincident with those of the direction of the LMC (Large Magellanic Cloud). Furthermore, we verified that all of these objects were included in the van Aarle et al. (2011) catalogue of LMC post-AGB stars, and 10 of them are also included in Kamath et al. (2015). We therefore decided to exclude these stars from our present study and focused on the remaining 157 Galactic post-AGB candidates.
@@
In Table LABEL:table:candidates we provide general data of the objects in this latter sample: a running number, the name, the Gaia DR3 ID, the J2000 coordinates, the G magnitude, the interstellar extinction in the V band, the spectral type, the reference for an identification as post-AGB (the Torun, Simbad, and Suárez et al. catalogues), and a flag providing information about binarity, variability, and the reason for excluding the source from the further analysis, as described in the previous section, together with a reference.
@@
The distance distribution of all these post-AGB candidates is shown in Fig. 2. We did not attempt to analyse the completeness of the sample since it presents several observational biases, including the impossibility of detecting the optical counterpart of some of the infrared sources with data in the IRAS catalogue.
3 Identification of Galactic post-asymptotic giant branch stars